1.    The fraction of defective integrated circuits produced in a photolithography process is being studied. A random sample of 300 circuits is tested, revealing 13 defectives.

a.    Calculate a 95% two-sided CI on the fraction of non-defective circuits produced by this particular tool.


b.    Calculate a 90% upper confidence bound on the fraction of defective circuits.


1
Expert's answer
2021-09-13T00:04:40-0400

Number of defective = 13

Total number of circuits = 300

a) Proportion of defective integrated circuits

"\hat{p} = \frac{x}{n}= \frac{13}{300}= 0.0433"

We will use Noramal distribution

"np= 300 \times 0.0433 = 13 \\

n(1-p) = 300 \times (1-0.0433) = 287"

95% two sided CI will be:

"\hat{p} -Z_{0.025} \sqrt{ \frac{\hat{p}(1- \hat{p})}{n}} ≤ p ≤ \hat{p} + Z_{0.025} \sqrt{ \frac{\hat{p}(1- \hat{p})}{n}} \\

Z_{0.025} = 1.96 \;(from \;table) \\

0.0433 - 1.96 \times 0.011 ≤ p ≤ 0.0433 + 1.96 \times 0.011 \\

0.02274 ≤ p ≤ 0.06486"

b) 95% upper confidence bound

"p ≤ \hat{p} + Z_{0.05} \sqrt{\frac{\hat{p}(1- \hat{p})}{n}} \\

Z_{0.05}= 1.645 \\

p ≤ 0.0433 + 1.945 \sqrt{\frac{0.0433 \times 0.9567}{300}} \\

p ≤ 0.0627"